An elevator weighing 2500 N ascends at a constant speed of 8.0 m/s. How much power must the motor supply to do this?
Answer #1
Power = work / time = fore* displacement/time = force* velocity
= 2500 * 8 = 20000 watts = 20 KW.
Answer #2
power=Force * velocity
=2500*8=2*10^4 W
Answer #3
P = force . velocity = 2500 x 8 = 20000 W/s